OpenStudy (e_s_j_f):

hey can somone please help me with this question. intergral =tan(x)^3*secx^(5) dx

OpenStudy (e_s_j_f):

its using trig identities

OpenStudy (turingtest):

notice that \[\frac d{dx} \sec x=\sec x\tan x\]strip that out and see what you have to work with

OpenStudy (turingtest):

in cases like these it is nice to shoot for even powers in the integrand so we can apply trig identities

OpenStudy (turingtest):

\[\int\tan^3x\sec^5xdx=\int\tan^2\sec^4\cdot\tan x\sec xdx\]this should give you some ideas

OpenStudy (e_s_j_f):

i have done that far. and differentiated it. but i dont know where to subsitute the trig identity.

OpenStudy (turingtest):

doesn't it look like \(\tan x\sec x\) will be our \(du\) ? If that is the case, what is our \(u\) ?

hartnn (hartnn):

try this alternate approach too :) write \(\Large \tan^2 x = \sec^2 x-1\) and then plug in u = sec x

hartnn (hartnn):

oh, Turing is going with that approach only!

OpenStudy (turingtest):

I was getting there @hartnn :P

OpenStudy (turingtest):

lol

OpenStudy (e_s_j_f):

\[\int\limits_{}^{}(\sec ^{2}x-1)\sec x^{4}.\tan x^{}\sec x^{} dx\]

OpenStudy (turingtest):

looking good, now simplify and u-sub

OpenStudy (e_s_j_f):

\[dx = du/secxtanx\]

OpenStudy (turingtest):

yes

OpenStudy (e_s_j_f):

do i cancel secxtanx

OpenStudy (turingtest):

I like to think of it like\[du=\sec x\tan xdx\]and then substituting for \(\sec x\tan x\), but your "cancellation" method will have the exact same result; it's just different logic

OpenStudy (e_s_j_f):

i see

OpenStudy (e_s_j_f):

so i have to integate u^5(u^2-1)

OpenStudy (turingtest):

yep

OpenStudy (turingtest):

er u^4(u^2-1)

OpenStudy (e_s_j_f):

I = secx^7/7 - secx^5/5 + C

OpenStudy (turingtest):

sounds correct, let's double check with the wolf

OpenStudy (anonymous):

do u stoll need help